# An Adversarial-Driven Experimental Study on Deep Learning for RF Fingerprinting

### v1

@@ Abstract
Index Terms:
Security, RF fingerprinting, Impersonation attacks, Deep learning
@@ I Introduction
  • To the best of our knowledge, we are the first to systematically evaluate DL-based RF fingerprinting systems from a security perspective. Unlike prior works that primarily focus on classification accuracy and domain robustness, we reveal a critical and underexplored vulnerability: the consistent misclassification behavior of CNN models under domain shifts can be exploited an effective backdoor to enable external attackers to launch impersonation attacks and intrude into the system.
@@ II-A Input/Output Modeling with RF Fingerprints
where $n=0,1,\dots,N-1$, $T_{s}$ is the sampling period of the OFDM system, and $w[n]$ is additive white Gaussian noise. $r[n]=\sum\limits_{l=0}^{L-1}h[l]x[n-l]$, the composite time-varying CSI $h[l]=p_{T}(nTs)\otimes g(nTs,lT_{s})\otimes p_{R}(nTs)$ includes the transmit filter $p_{T}(t)$, the receive filter $p_{R}(t)$, and the time-varying impulse response of the physical fading channel $g(t,\tau)$. The composite CFO is $\Delta\omega_{RT}=\Delta\omega_{R}-\Delta\omega_{T}$, and the composite phase offset is $\psi_{RT}=\psi_{R}-\psi_{T}$. The parameters $\eta=(\alpha_{R}\alpha_{T}+\beta_{R}\beta_{T}^{*})$, $\kappa=(\alpha_{R}\beta_{T}+\beta_{R}\alpha_{T}^{*})$, where $\alpha_{(\cdot)}$ and $\beta_{(\cdot)}$ represent the effects of IQ imbalance of either the transmitter or the receiver. W.l.o.g, for the transmitter $\alpha_{T}=\cos(\Delta\phi_{T})+j\epsilon_{T}\sin(\Delta\phi_{T})$ and $\beta_{T}=\epsilon_{T}\cos(\Delta\phi_{T})-j\sin(\Delta\phi_{T})$, where $\epsilon_{T}$ and $\Delta\phi_{T}$ represent the amplitude and phase differences between the transmitter’s IQ branches.
@@ IV In-Lab Experimental Setup
$\bullet$ TrS 2: Collected on Day 2 at time $t_{3}$, with TXs sending repeated signals at location $l_{4}$.
@@ V-A Classifier Architecture
In this work, we adopt the CNN architecture presented in [5], which is a typical design consisting of eight layers including four convolutional layers followed by three fully connected (dense) layers. All input data are normalized using per-device standardization, where for each deice, we flatten its I/Q data and fit using only that device’s data before being fed into the CNN model. For the input, a sliding window of frame size 4 is implemented that makes the input as 2$\times$256 where the I and Q components of the signal are stored in two rows of matrix. Each of the convolutional layers uses 40 filters, where the first two have kernel sizes of 1$\times$7 and 1$\times$5, respectively, each followed by a max-pooling layer that reduces the dimensionality to 2$\times$128$\times$40 and 2$\times$64$\times$40, respectively. These are followed by a convolution layer of size 2$\times$7 and max-pooling layer that further reduces the dimension to 2$\times$32$\times$40. Finally, a convolution layer of size 2$\times$5 is implemented followed by a flattening layer, and the final feature maps are passed through three dense layers with 1024, 256 and 4 neurons, respectively. Dropout is applied after the flattening and first dense to help prevent overfitting. The final classification layer uses softmax activation to output probabilities over 4 classes. We use Adam optimizer with a learning rate of 1×$\times 10^{-3}$ to train the model.
@@ V-B1 Consistent misclassification behavior during domain shifts
To verify whether this misclassification behavior of the model during domain shifts is consistent and allows Eve to easily penetrate the system by impersonating a legitimate transmitter, we show the success rate of the replay attack launched by Eve. In this untargeted attack, Eve attempts to impersonate any legitimate device by replaying AS 1. As illustrated in Fig. 3(b), we observe that Eve can successfully impersonate Devices 2, 3, and 4 with relatively high probabilities ($\geq$ 95%) and an overall untargeted impersonation attack success rate of 95.96%.

### v2

# Xinyu Cao11  Bimal Adhikari12  Shangqing Zhao1  Jingxian Wu2  Yanjun Pan2 1Co-First Authors

@@ I Introduction
  • To the best of our knowledge, we are the first to systematically evaluate DL-based RF fingerprinting systems from a security perspective. Unlike prior works that primarily focus on classification accuracy and domain robustness, we reveal a critical and underexplored vulnerability: the consistent misclassification behavior of CNN models under domain shifts can be exploited as an effective backdoor to enable external attackers to launch impersonation attacks and intrude into the system.
@@ II-A Input/Output Modeling with RF Fingerprints
where $n=0,1,\dots,N-1$, $T_{s}$ is the OFDM sampling period of the OFDM system, and $w[n]$ is additive white Gaussian noise. $r[n]=\sum\limits_{l=0}^{L-1}h[l]x[n-l]$, the composite time-varying CSI $h[l]=p_{T}(nTs)\otimes g(nTs,lT_{s})\otimes p_{R}(nTs)$ includes the transmit filter $p_{T}(t)$, the receive filter $p_{R}(t)$, and the time-varying impulse response of the physical fading channel $g(t,\tau)$. The composite CFO is $\Delta\omega_{RT}=\Delta\omega_{R}-\Delta\omega_{T}$, and the composite phase offset is $\psi_{RT}=\psi_{R}-\psi_{T}$. The parameters $\eta=(\alpha_{R}\alpha_{T}+\beta_{R}\beta_{T}^{*})$, $\kappa=(\alpha_{R}\beta_{T}+\beta_{R}\alpha_{T}^{*})$, where $\alpha_{(\cdot)}$ and $\beta_{(\cdot)}$ represent the effects of IQ imbalance of either the transmitter or the receiver. W.l.o.g, for the transmitter $\alpha_{T}=\cos(\Delta\phi_{T})+j\epsilon_{T}\sin(\Delta\phi_{T})$ and $\beta_{T}=\epsilon_{T}\cos(\Delta\phi_{T})-j\sin(\Delta\phi_{T})$, where $\epsilon_{T}$ and $\Delta\phi_{T}$ represent the amplitude and phase differences between the transmitter’s IQ branches.
@@ IV In-Lab Experimental Setup
$\bullet$ TrS 2: Collected on Day 2 at time $t_{3}$, with TXs sending repeated signals at location $l_{3}$.
@@ V-A Classifier Architecture
[…]
@@ V-B1 Consistent misclassification behavior during domain shifts
To verify whether this misclassification behavior of the model during domain shifts is consistent and allows Eve to easily penetrate the system by impersonating a legitimate transmitter, we show the success rate of the replay attack launched by Eve. In this untargeted attack, Eve replays AS 1 to impersonate any legitimate device. As illustrated in Fig. 3(b), we observe that Eve can successfully impersonate Devices 2, 3, and 4 with relatively high probabilities ($\geq$ 95%) and an overall untargeted impersonation attack success rate of 95.96%.